In 2018 Lamborghini was quite concerned with pricing the new Lambo, the Lambo Centenario. The manager is interested in maximizing revenue. The demand schedule for the Lambo Centenario is given by P=7 – (1/10) Q. The company plans to sell the Lambo Centenario for about $2M. (a) How many lambos is the company planning to sell? (b) What is the elasticity of demand at this given price? (c) Is this the price that maximizes revenue?
Answer #1

Demand is P = 7 - Q/10. This implies revenue is PQ or R = 7Q - Q^2/10 and marginal revenue, that has twice the slope of demand with same intercept, is MR = 7 - 2Q/10. When MR is 0, the firm is maximizing revenues. Hence revenue maximizing quantity is Q = 7*10/2 = 35 and revenue maximizing price is P = 7 - 35/10 = $3.5 M

a) Company sells Q = (7 - 2)*10 = 50 lambos at this price.

b) We know that MR = P(1 - 1/e). Here P = 2 and MR = 7 - 2*50/10 = -3. This gives -3 = 2(1 - 1/e) or 1/e = 1 + 1.5. Thus, elasticity of demand ed = 0.4.

c) No. Revenue maximizing price is $3.5 M. Also, when price is $2 M, marginal revenue is negative and not zero.